OpenStudy (anonymous):

i will medal and fan let a,b,c be points (2,3), (3,-2) and (-1,4) find length BC and AC

OpenStudy (anonymous):

well first, you have to know the distance formula

OpenStudy (anonymous):

try to put theses on a graph

OpenStudy (anonymous):

then measure by cubic units from 1 to the other 1

OpenStudy (anonymous):

you will probally make a triangle so i prefer to get some paper out :P

OpenStudy (anonymous):

thanks sir. please am watching

OpenStudy (anonymous):

\[d=\sqrt{(x _{2}-x _{1})^{2}+(y _{2}-y_{1})^{2}}\]

OpenStudy (anonymous):

just plug the coordinates in

OpenStudy (anonymous):

options given are a)

OpenStudy (anonymous):

\[\sqrt{30},\sqrt{42},\sqrt{25},\sqrt{52} for BCplease help me. \]

OpenStudy (anonymous):

www.analyzemath.com/practice_tests/sat.../level_1_sample_1_sol.html this will help, read it